OpenStudy (anonymous):

1( u-v/8v) + (6u - 3v/ 8v) 2)(m - 3n/ 6m^3 n) - (m + 3n/6m^3 n) 3) (6/ x - 1) - (5x /4) 4. (3/ x + 7) + (4/ x - 8) 5. (7/3) - (8 / 12x - 8) 6. (2x/ 5x -4) + (6x/ 2x +3)

OpenStudy (primeralph):

What are we to do?

OpenStudy (anonymous):

Add or Subtract. Show work for problems you have to find the LCD.

OpenStudy (primeralph):

@uri Please take care of these.

OpenStudy (ivettef365):

on the first problem the LCD = 8V

OpenStudy (uri):

I think she posted like 6 Questions together. Hey,@_1kjazz,Welcome to Openstudy,Please post on Question at a time,Edit it and remove the other 5,1 question 1 thread,thanks.

OpenStudy (ivettef365):

so you have u-v 6u - 3v ----- + ------ 8V 8V

OpenStudy (primeralph):

@uri Please let it slide. I think she needs them urgently.

OpenStudy (anonymous):

@primeralph i do! @uri im sorry im kinda new to this @ivettef365 yes that is correct

OpenStudy (primeralph):

@ganeshie8 Could you help here? I got to go.

OpenStudy (ivettef365):

ok you add as if it was a common fraction so you have u - v + 6u - 3v --------------- 8v

OpenStudy (ivettef365):

now, combine like terms so you have 7u -4v ------- 8v

OpenStudy (ivettef365):

the next problem is similar, since you have a common denominator can you do it by yourself ??

OpenStudy (ivettef365):

the only difference now is that you are subtracting instead of adding

OpenStudy (anonymous):

@ivettef365 yes i can do it im still confused on the other ones though

OpenStudy (ivettef365):

ok, do #2, tell me what you got for answer and I will explain #3 in the meantime

OpenStudy (ivettef365):

you have the following 6 5x 6(4) -5x(x-1) ----- - ------ = ------------- x-1 4 4(x-1)

OpenStudy (ivettef365):

common denominator would be 4(x-1) do you understand what I did

OpenStudy (anonymous):

worked on number 2 i got -6/6m^3n i dont think its correct though

OpenStudy (ivettef365):

#2 m - 3n m + 3n m - 3n -m - 3n -------- - ---------- = ----------------- 6m^3n 6m^3n 6m^3n

OpenStudy (ivettef365):

yes, you got it right

OpenStudy (ivettef365):

but it should be -6n

OpenStudy (ivettef365):

-6n/6m^3n

OpenStudy (ivettef365):

and you can simplify even more you have -6n -1 ------- = ---------- 6m^3n m^3

OpenStudy (anonymous):

i think im getting it could we move on to number 4

OpenStudy (ivettef365):

sure, #4 would be set up like this 3 4 3(x-8) + 4(x+7) ------- + ------- = _________________________ x+7 x-8 (x+7)(x-8)

OpenStudy (anonymous):

would that be the final answer

OpenStudy (ivettef365):

I meant 3x - 24 + 4x + 28 ---------------- (x+7)(x-8)

OpenStudy (anonymous):

then just solve from there ?

OpenStudy (ivettef365):

right, so you have 7x + 4 --------- there is nothing else you can do on this one (x+7)(x-8)

OpenStudy (anonymous):

cool beans could we do i now

OpenStudy (anonymous):

5

OpenStudy (ivettef365):

now you have 7 8 7(12x-8) + 8(3) ------- - -------- = ---------------- 3 12x-8 3(12x-8)

OpenStudy (ivettef365):

now you have 84x -56 - 24 ------------- 36x-24

OpenStudy (ivettef365):

do you see what I did, on the top should be 7(12x-8) - 8 (3)

OpenStudy (ivettef365):

I put + by mistake, do you see it

OpenStudy (anonymous):

yes i do

OpenStudy (ivettef365):

ok now you have 84x - 80 -------- 36x -24

OpenStudy (ivettef365):

now you can simplify even more, you have 4(21x - 20) 21x-20 ---------- = ---------- 4(9x - 6) 9x - 6

OpenStudy (anonymous):

alright ithink im getting this

OpenStudy (ivettef365):

ok, try to do 6

OpenStudy (anonymous):

2x(2x+3) 6x(5x-4) ___________ + _____________ 5x-4 2x+3

OpenStudy (anonymous):

2x(2x+3)+6x(5x-4) --------------------- (5x-4)(2x+3)

OpenStudy (anonymous):

4x^2+6x+30x^2-24x ______________________ (5x-4)(2x+3)

OpenStudy (anonymous):

-18x + 34x2 ______________________ (5x-4)(2x+3)

OpenStudy (anonymous):

is this the correct final answer ?

OpenStudy (anonymous):

@ivettef365 ?

OpenStudy (ivettef365):

Im doing it, give me a minute

OpenStudy (ivettef365):

that is what I get

OpenStudy (ivettef365):

let me see if it could be simplify more

OpenStudy (anonymous):

okay :)

OpenStudy (ivettef365):

no, it can not be simplify any further, that is good

OpenStudy (anonymous):

thanks would you be willing to help me with multiplication ?

OpenStudy (ivettef365):

sure

OpenStudy (anonymous):

1. (x + 3)/4) * 3(x -6)/ 3(x + 3) 2. (x + 8)/(x +6)(x -8) * 4x(x + 10)/( x + 10) 3. (28 - 7b)/ (b -4) * (1/ ( b + 10) 4. (1 / n+ 9) * (3n - 18)/ ( 6 - n) 5. ( 2b2 - 12b)/ ( b + 5) * ( b + 5)/ ( b - 6)

OpenStudy (ivettef365):

ok, multiplication is easier than adding so you have x + 3 3(x-6) ---------- x ----------- 4 3(x + 3)

OpenStudy (ivettef365):

now do you see what numbers you can cancel, whatever is the same on the top and the bottom you can cancel, therefore you have x + 3, on top and bottom, and you have a "3" on top by itself and a "3" on bottom by itself do you see it

OpenStudy (anonymous):

yes i think im starting to get it

OpenStudy (ivettef365):

ok so when you eliminate from top and bottom then you have 1 x-6 ------- x ---------- 4 1

OpenStudy (ivettef365):

you see I put a 1 on the numbers eliminated the reason why is because when you have 3 --- = 1 3

OpenStudy (ivettef365):

so the same number on top and bottom is going to be = 1

OpenStudy (ivettef365):

so now you can multiply and you have x - 6 -------- 4

OpenStudy (anonymous):

final answer ?

OpenStudy (ivettef365):

yes

OpenStudy (anonymous):

2 ?

OpenStudy (ivettef365):

are you sure problem 2 is written correctly ??

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

ok, lets skip that to the end, the others are easy, so lets do #3

OpenStudy (ivettef365):

you have 28 - 7b 1 ------- x -------- b - 4 b+10

OpenStudy (ivettef365):

you can write this problem as follows: -7(b-4) 1 -------- x -------- (do you see how I convert 28 - 7b) b-4 b+10

OpenStudy (anonymous):

yes it looks like you broke it down

OpenStudy (ivettef365):

right that is the same when you multiply -7(b-4) = -7b+28 = 28 -7b

OpenStudy (ivettef365):

now which numbers can you elimiate

OpenStudy (anonymous):

b-4,cause there the same right ?

OpenStudy (ivettef365):

right so what do you get

OpenStudy (anonymous):

well i didnt do much after that all i knew was to cross out b-4

OpenStudy (ivettef365):

ok, so you cross b-4, now what do you get when you multiply ??

OpenStudy (ivettef365):

you have left the following -7 1 ------ x ------- = ?? 1 b+10

OpenStudy (anonymous):

will i have -7/b+10

OpenStudy (ivettef365):

good job

OpenStudy (anonymous):

could we go to the next one ?

OpenStudy (ivettef365):

ok, the next one you have 1 3n-18 ------ x ---------- n+9 6-n

OpenStudy (ivettef365):

now 3n -18 = -3(6-n) (do you see it)

OpenStudy (anonymous):

um no im not understanding

OpenStudy (ivettef365):

how would you multiply -3(6-n)

OpenStudy (anonymous):

i would distribute it would be -18+3n

OpenStudy (ivettef365):

good, so that is the same as 3n-18 right

OpenStudy (ivettef365):

now you have 1 -3(6-n) ------- x ---------- n+9 6-n

OpenStudy (ivettef365):

what can you eliminate ??

OpenStudy (anonymous):

6-n

OpenStudy (ivettef365):

ok so what do you have left

OpenStudy (anonymous):

1 -3(1) _____ * _____ n+9 1

OpenStudy (ivettef365):

good job, now what is the answer

OpenStudy (anonymous):

-3 _____ n+9

OpenStudy (ivettef365):

very good, you want to try doing the last one

OpenStudy (anonymous):

umm sure

OpenStudy (anonymous):

( 2b2 - 12b) ( b + 5) ____________ *_________ = (b+5) (b-6)

OpenStudy (anonymous):

( 2b2 - 12b) 1 ____________ *_________ = 1 (b-6)